I have tried to work this problem out several times.
2x+3y=5
x=3y+1
Using substition I have tried 2x+3(1-x)=5
2x+3-3x=5
-3 -3
________
2x-3x =2
-1x =2
x=-2 but this solution does not work

Found 2 solutions by sarah_adam, checkley71:
Answer by sarah_adam(201)

You substituted the wrong value of x
given
2x+3y=5
x=3y+1
substitute value of x in the equation 2x+3y = 5
2(3y+1)+3y = 5
6y+2+3y = 5
on solving we get 9y = 3
therefore y = 3/9 = 1/3
therefore x = 3(1/3) + 1 = 2
therefore x = 2 and y = 1/3

Answer by checkley71(8403)

2x+3y=5
x=3y+1 substituting (3y+1) for x in equation #1 we get
2(3y+1)+3y=5 or 6y+2+3y=5 or 9y=3 or y=3/9 or y=1/3
substituting 1/3 for y in the second equation we get
x=3(1/3)+1 or x=1+1 or x=2
proof 2*2+3(1/3)=5 or 4+1=5 or 5=5 and
2=3(1/3)+1 or 2=1+1 or 2=2